OpenStudy (anonymous):

Which function below represents the arithmetic sequence 3, 7, 11, 15...? A)f(n)=4+3(n-1) B)f(n)=4+3n C)f(n)=3+4n D)f(n)3+4(n-1)

OpenStudy (asib1214):

C. 3+4*0=3 3+4*1=7 3+4*2=11 3+4*3=15